help please:
“Divide 45 into two parts so that if 4 is subtracted from the lager number the result is 7 more than the smaller number?”


My instinct tells it is something to do with subtracting 11 from 45 and finding a prime number – but I am splashing about in a muddy puddle me thinks – what is the equation I should be using
how do I do this ?

Answer by RajM(5)

We set up equations first lets assume larger number is x and smaller number is y.
x+y=45
x-4=y+7
x-y=11
x+y=45
x-y=11
we use elimination method hence we add both equations and eliminate y
2x=56
x=56/2
x=28
next we plug x in any equation and solve for y .
x+y=45
28+y=45
y=17
Hence numbers are 28 and 17.